QUESTION 40

  1. Amount of Variable Input

    Amount of Output

    Unit Price

    Cost per input

    Total cost

    Marginal cost

    0

    0

    0

    0

    0

    0

    1

    5

    $45

    $670

    2

    25

    $45

    $670

    3

    55

    $45

    $670

    4

    75

    $45

    $670

    A

    5

    100

    $45

    $670

    B

    6

    118

    $45

    $670

    C

    F

    7

    135

    $45

    $670

    D

    G

    8

    152

    $45

    $670

    E

    H

    9

    167

    $45

    $670

    I

    10

    170

    $45

    $670

    J

    11

    169

    $45

    $670

    What is A+B+C+D+E?

  2. What is F+G+H+I+J? (Use table in question 40) ignore decimal places

  3. Using results of question 37-40, find profit maximizing output level.

    From output level between  and
Answer #1

Total cost = Amount of variable input*Cost per input
A = 4*670 = 2,680
B = 5*670 = 3,350
C = 6*670 = 4,020
D = 7*670 = 4,690
E = 8*670 = 5,360

So,  A+B+C+D+E = $20,100

Marginal cost = Change in total cost/Change in output
F = (C-B)/(118-100) = (4,020-3350)/18 = 670/18 = 37.22
G = (D-C)/(135-118) = (4,690-4,020)/17 = 670/17 = 39.41
H = (E-D)/(152-135) = (5,360-4,690)/17 = 670/17 = 39.41
I = 670/(167-152) = 670/15 = 44.67
J = 670/(170-167) = 670/3 = 223.33
(Change in total cost = 670)

F+G+H+I+J = $384.04

Profit maximizing output level is that for which MR just exceeds or equals MC.
MR = 45 exceeds MC = 44.67 till Q = 9. After Q = 9, MR < MC.
So, the profit maximizing output level is 9 units.